f(0)=f(1)=0,f / (1)=2 &y=(ex   )ef(x)
dy/dx=
 
 

Grade:12

1 Answers

vivek
8 years ago
let,
   g’(0)=f’(1)ef(0)+f(1)ef(0).f(0)
=f’(1)+f’(0)=0+2=2